# Molecular clouds: do they deserve a non-Gaussian description?

### v1

@@ 4.1 One-point based statistics
We list below the one-point statistics used in this paper. Even though some of these statistics can probe non Gaussian information such as sparsity, they are all point wise statistics. This means in particular that they cannot capture spatial arrangement in the maps.
@@
  • The quantile functions [^6] We consider quantile functions rather than probability distribution functions, as the difficulty of defining a unique binning range makes comparisons less efficient. Indeed, to define a range in terms of quantiles allows for a natural adaptation to each MCs’ regions. (QF) normalized by the median:
@@
    where $q_{\alpha}(x)$ designates the $\alpha$-quantile of the distribution of values $\{x(\@vec{u})\}_{\@vec{u}}$. Hence $q_{1/2}$ stands for the median operator. We consider 10 quantiles $\{\alpha_{i}\}_{i}$ such that $1-\alpha_{i}$ are logarithmically spaced between $10^{-4}$ and $0.4$. Low quantiles are not considered as, for observational data, they are contaminated (Ossenkopf-Okada et al., 2016). This logarithmic binning of the high column density values is motivated by the log-normal to power-law behavior of the tail of MCs’ PDFs. See for instance Pouteau et al. (2023) and references therein.
@@ 4.2 Two-point based statistics
A very popular way to describe spatial properties of a process is through the power spectrum (PS). It is defined as the Fourier transform of the auto-correlation function, and describes the energy distribution scale by scale in the process studied. In this paper, we consider the isotropic PS, defined as:
@@
A power-law behavior $PS[k]\propto k^{-\beta}$ is expected for fields arising in turbulent MHD (Schekochihin, 2022). Thus we adopt a log-log representation, to linearize it and for stability purposes (Bruna & Mallat, 2013):
@@
where $x_{\text{apo}}$ is apodized to mitigate non periodic boundary conditions (PBC) as explained in appendix B. We use 6 band-pass filters $\{\psi_{i}\}_{1\leq i\leq 6}$ defined as:
$\tilde{\psi}_{i}[\@vec{k}]=\mathrm{1}_{\|\@vec{k}\|\in[k_{i},k_{i+1}[},$
@@ 4.3 Scattering statistics
The Wavelet Scattering Transform (WST) is a set of non-Gaussian descriptors with a hierarchical and multi-scale structure (Bruna & Mallat, 2013). It has been shown to be very efficient at describing astrophysical fields (see, for instance, Allys et al., 2019; Saydjari et al., 2021). The usual WST consists in two layers of statistics, the first of which depends on a single scale of length $\simeq 2^{j}$ pixels with orientation $\theta$:
@@
    Note that it computes a $L^{1}$ norm of the filtered field, in contrast with the power spectrum that probes a $L^{2}$ norm.
  • $S_{2}^{Iso1}$ and $S_{2}^{Iso2}$, that measure an isotropic coupling between scales and are defined by fitting the following model:
    $\log_{2}\,S_{2}(x)[j_{1},j_{2},\theta_{1},\theta_{2}]\simeq S_{2}^{Iso1}(x)[j_{1},j_{2}]\,+\,\\ S_{2}^{Iso2}(x)[j_{1},j_{2}]\cos\big{[}2(\theta_{2}-\theta_{1})\big{]}.$
@@
In this paper, we will furthermore compress both $S_{2}^{Iso1}$ and $S_{2}^{Iso2}$ coefficients that depend on two scales $(2^{j_{1}},2^{j_{2}})$ by only keeping their $j_{2}-j_{1}$ dependency, i.e., a dependency on their ratio, by considering the following average:
@@
In practice, we consider 4 scales between $j_{\min}=2$ and $j_{\max}=5$, which leads to the 3 possible values $1\leq\delta\leq 3$ for our reduced $S_{2}$ coefficients. We also divide the half-plane $[0,\pi[$ in 8 different $\theta$ angles. For the finest maps of pixel size $6\arcsec$, this $j_{\min}$ corresponds to a scale of $24\arcsec$ that remains above the $18\arcsec$ resolution of the observations.
For RWST computations, convolutions are done with local wavelets. This allows us to mitigate non PBC after convolution by suitably cropping the corrupted edges, without using apodization. These computations are done with the pywst [^7] https://github.com/bregaldo/pywst Python package (Regaldo-Saint Blancard et al., 2020).
@@
Gaussian description interactions
@@
Table 2: Overview of the sets of summary statistics, their dimensions and properties (whether they are sensitive to non-Gaussianity, provide a multiscale description, and probe couplings between scales).
4.4 Overview
The statistics introduced above are compiled in Table 2. In the following, in addition to the sets of summary statistics defined above, we will also consider some aggregated sets of summary statistics that are made from groups of these building blocks. For instance, we will refer to $\phi_{\text{Gaussian}}$ statistics to mean the joint set of $\{\phi_{\text{mean}},\phi_{\text{var}},\phi_{\text{PS}}\}$. These groups are defined in Table 3.
@@ 4.4 Overview
final log-Gaussian, RWST 17
Table 3: Aggregated sets of summary statistics used in this work.
5 Towards a low-degeneracy set of statistics
We want to construct a set of informative summary statistics for the observational dataset. To do so, we will follow a bottom-up approach, starting with usual low-order statistics, and increasingly trying to improve on them, by exhibiting and lifting potential degeneracies. The underlying idea is that low-order statistics are good candidates to concentrate most of the informative power into a few coefficients. Such a concentration property is of interest, especially as our low data regime prevents us from learning many features. Another benefit of this approach is to first promote simple statistics, and use more elaborate ones only if they add a significant contribution.
@@ 5.1 Molecular clouds have Gaussian degeneracies
To begin with, we investigate whether Gaussian statistics (i.e. the set made of mean and PS statistics) are degenerate for the observational data. To point out such potential degeneracies, we confront these Gaussian statistics, starting with a low-order non-Gaussian set of statistics, the quantile function QF, using the methodology introduced in Sec. 3.3. As shown in Fig. 7, this first confrontation on observational data already underlines a strong degeneracy level for both Gaussian and QF features, showing that neither statistics is by itself sufficient, according to the compatibility diagnostic we introduced. In particular, the quantile statistics extract a significant amount of information that cannot be efficiently captured by Gaussian statistics.
@@
Figure 7: Confronting Gaussian statistics with QF statistics on observational data, based on the test presented in Fig. 6. Each set of statistics has strong degeneracies lifted by the other set. To investigate the Gaussian confusions, we pick six pairs of 512 $\times$ 512 patches, whose locations on the scatter plot are given by the red stars. These pairs are shown in Fig. 8.
@@
Figure 8: Examples of Gaussian confusions. Six pairs of 512 $\times$ 512 patches are chosen, whose locations on the scatter plot of Fig. 7 are given there by the red stars. The column density maps are shown in units of $10^{20}\text{cm}^{-2}$. For each patch, we report: $s$ the sub-sampling factor from the original 3″/pix map, $d$ and $b$ the approximated distance and Galactic latitude of the cloud. The pixel size (in  mpc) of a patch is thus proportional to $s\times d$. If a pair has patches $(i,j)$ with incompatible pixel sizes, that we define according to the following criterion $\max\{\frac{s_{i}d_{i}}{s_{j}d_{j}},\frac{s_{j}d_{j}}{s_{i}d_{i}}\}\geq 3/2$, we color the labels in red.
@@
Figure 9: Statistics for the examples of Gaussian confusions shown in Fig. 8. In each row, the orange filled line (resp. band) corresponds to the mean (resp. std) of the statistics computed over the four 256 $\times$ 256 sub-patches of the top patch of each pair of Fig. 8, and the corresponding blue lines and areas refer to the bottom patch of the pair. The top row corresponds to the statistics used in the $x$-axis of the scatter plot of Fig. 7, i.e., QF statistics, plotted with 10 increasing quantile values, while the bottom row corresponds to the $y$-axis, i.e., Gaussian statistics, plotted starting with mean and followed by the binned PS with six decreasing scales. To better highlight the discrepancies between the two patches of a given pair, we report in second and third rows the offsets of the orange and blue filled lines with respect to their common mean.
@@
In addition to this dataset-wide diagnostic, we display in Fig. 8 six randomly picked pairs of 512 $\times$ 512 patches degenerate for Gaussian statistics but with increasing QF incompatibility. The locations of these pairs on the scatter plot of Fig. 7 are given there by the red stars. For these six pairs of patches, we show in Fig. 9 the two sets of statistics used in the diagnostic (QF statistics and Gaussian statistics), where the orange filled lines (resp. bands) correspond to the mean (resp. standard deviation) of the statistics computed over the four 256 $\times$ 256 sub-patches of the top patch of each pair of Fig. 8, and the corresponding blue lines and areas refer to the bottom patch of the pair. To better highlight the discrepancies between the two patches of a given pair, the offsets of the orange and blue filled lines with respect to their common mean are also shown in Fig. 9. These results illustrate the ability of QF statistics to distinguish between images with the same Gaussian statistics, which confirms that our compatibility diagnostic works as expected. We emphasize here that the $\phi_{\text{QF}}$ statistics are complementary to the one-point properties probed by Gaussian statistics.
@@
This behavior is not particularly surprising as one-point properties of column density maps of MCs are expected to be at least log-normal, if not with a power-law tail. Hence, estimating such properties directly from the logarithm of the maps, as does $\phi_{\text{QF}}$ but not $\phi_{\text{Gaussian}}$, might enhance the discriminative power. To test this idea, we study in the first column of Fig. 10 the degeneracy level of the set of mean and variance statistics estimated respectively on the raw maps and their logarithm. More precisely, top row plots confront QF statistics to $\{\phi_{\text{mean}},\,\phi_{\text{var}}\}$, while bottom row plots confront QF statistics to $\{\phi_{\text{mean of log}},\,\phi_{\text{var of log}}\}$. We see in this figure that the pairs of observations are clearly better discriminated when using the mean and variance computed on the logarithm of the maps (d) rather than directly (a). This effect is very well reproduced by logFBM data (middle column), but does not hold in general, as shown with the Describable Textures Dataset (DTD) (right column).
@@
Surprisingly, this analysis also shows that, for datasets such as observations or logFBMs, probing the PDF properties through the prism of a mean and a variance is more discriminative than probing its shape. Here, in both d) and e), our suitably constructed set $\{\phi_{\text{mean of log}},\,\phi_{\text{var of log}}\}$, of dimension 2, performs almost always a better discrimination than the set $\phi_{\text{QF}}$, of dimension 10. This emphasizes the importance of suitably constructed low dimensional descriptions in such analysis. The performative power of such low-dimensional features is usually due to the exploitation of underlying regularities of the studied processes that allow these compression, without much loss of information. For deeper insights on the connection between regularity, approximation and sparsity, we recommend Pr. Mallat’s lecture [^8] https://www.college-de-france.fr/fr/agenda/cours/representations-parcimonieuses/le-triangle-regularite-approximation-parcimonie. However, when dealing with datasets made of a wide diversity of irregular processes, such as everyday life textures like DTD, it is hard to identify a compression that does not lead, for some pairs, to poorer performances (c, f).
Figure 10: Top row plots confront QF statistics to $\{\phi_{\text{mean}},\,\phi_{\text{var}}\}$, while bottom row plots confront QF statistics to $\{\phi_{\text{mean of log}},\,\phi_{\text{var of log}}\}$. These results evidence that taking the logarithm of the map enhances the discriminative power of the mean and variance statistics on both observation and logFBM data (left and middle columns) but not on DTD (right column).
@@ 5.2 Molecular clouds have log-Gaussian degeneracies
The ability of Gaussian one-point statistics to grasp efficiently one-point properties of observations from the logarithm of the column density maps, shown in Fig.10, suggests to shift towards log-Gaussian statistics, i.e., mean and PS estimated on the logarithms of the maps. We thus investigate now whether we may point out some degeneracies of these statistics on observational data. However, if we confront this set with QF statistics, as we did previously for Gaussian statistics, we do not expect to lift significant log-Gaussian degeneracies. Instead, we suggest to search for such degeneracies using a higher order set: the RWST. To make sure that, on other datasets, the potential degeneracies lifted by these higher order statistics could not be lifted by one-point statistics, we will confront the log-Gaussian set to the RWST one.
@@
We show in Fig. 11 that log-Gaussian statistics have degeneracies on observational data lifted by RWST. We show six examples of degenerate pairs in Fig. 12 and the statistics of these patches in Fig. 13. In most of these pairs, the $S_{1}^{Iso}$ coefficients heavily contribute to lift the log-Gaussian degeneracy. In the first and last pairs, $\langle S_{2}^{Iso2}\rangle_{j_{2}-j_{1}}[\delta]$ and respectively $\langle S_{2}^{Iso1}\rangle_{j_{2}-j_{1}}[\delta]$ are also very discriminative.
@@
When applying the same analysis on logFBM data (Fig. 14.a), this finding however does not hold. The RWST diagnostic brings no additional information that was not already probed by log-Gaussian statistics. Indeed, as expected for such a dataset, log-Gaussian statistics are sufficient (Cover, 1999). Finally, simulations yield yet another result (Fig. 14.b). Although this latter case leads to the same qualitative conclusions as observations, namely underlining the insufficiency of log-Gaussian statistics, it differs from it quantitatively. This supports the caveat of simulation-based inference according to which observations and simulations of such processes lie in sub-manifolds having different geometries, and motivates the observation-based approach of this paper.
5.3 Final set of statistics
In this paper, we choose to limit our study to the log-Gaussian + RWST set, which we call $\phi_{\text{final}}$, that seems well suited to describe this observational dataset. To ensure this suitability any further, we could continue to study the degeneracies of $\phi_{\text{final}}$ by confronting it with other sets of statistics. In any case, it is difficult to guarantee with certainty at any stage that we have not missed any remaining degeneracy of the resulting set that would be obtained by such a construction, even if each confrontation with a complementary set reinforces our confidence in its completeness.
Once a set of statistics has been fixed, however, it is also possible to use a complementary diagnostic to study it, which consists in checking the visual proximity of maps identified as close. If they are clearly visually different, this indicates complementary limitations of the set of statistics used. This is done with $\phi_{\text{final}}$ in the following section, where we introduce a distance to evaluate a notion of proximity rather than compatibility between pairs of maps.
Finally, we want to emphasize that low-order statistics generally constitute a precious and easily-accessed source of information. They should not be underestimated, and attempting to tailor them to the type of data considered can be fruitful, as shown here when shifting from a Gaussian to a log-Gaussian description.
Figure 11: Same as Fig. 7, but using the log-Gaussian and RWST sets of statistics on the observational dataset. To investigate the degeneracies of the log-Gaussian statistics, six pairs of patches are selected, corresponding to the six red stars on this plot.
Figure 12: Examples of log-Gaussian degeneracies. Six pairs of 512 $\times$ 512 patches are chosen, whose locations on the scatter plot of Fig. 11 are given there by the red stars. The column density maps are shown in units of $10^{20}\text{cm}^{-2}$. For each patch, we report: $s$ the sub-sampling factor from the original 3″/pix map, $d$ and $b$ the approximated distance and Galactic latitude of the cloud. If a pair has patches $(i,j)$ with incompatible pixel sizes according to the following criterion $\max\{\frac{s_{i}d_{i}}{s_{j}d_{j}},\frac{s_{j}d_{j}}{s_{i}d_{i}}\}\geq 3/2$, we color the labels in red.
Figure 13: Statistics for the examples of log-Gaussian confusions shown in Fig. 12. In each row, the orange filled line (resp. band) corresponds to the mean (resp. std) of the statistics computed over the four 256 $\times$ 256 sub-patches of the top patch of each pair of Fig. 12, and the corresponding blue lines and areas refer to the bottom patch of the pair. The top row corresponds to the RWST, starting with $S_{1}^{Iso}[j]$ coefficients with four increasing scales $j$, then $\langle S_{2}^{Iso1}\rangle_{j_{2}-j_{1}}[\delta]$ coefficients with three increasing scale ratios $\delta$ and finally $\langle S_{2}^{Iso2}\rangle_{j_{2}-j_{1}}[\delta]$ coefficients with the same three scale ratios. The bottom row represents in the following order: mean of log followed by PS of log with 6 decreasing scales. The second and third rows show the offsets of these statistics with respect to the mean of the two.
Figure 14: log-Gaussian statistics have degeneracies lifted by RWST for simulations in (b), but not for logFBM processes in a). This is expected because log-Gaussian statistics are sufficient for families made of logFBM processes (Cover, 1999).

### v2

@@ 4.1 One-point based statistics
We list below the one-point statistics that are used in this paper. Even though some of these statistics can probe non-Gaussian information such as sparsity, they are all pointwise statistics. This means in particular that they cannot capture spatial arrangement in the maps.
@@
  • The quantile functions [^8] We consider quantile functions rather than probability distribution functions, as the difficulty of defining a unique binning range makes comparisons less efficient. Indeed, to define a range in terms of quantiles allows us for a natural adaptation to each MCs’ regions. (QF) normalized by the median:
@@
    where $q_{\alpha}(x)$ designates the $\alpha$-quantile of the distribution of values $\{x(\@vec{u})\}_{\@vec{u}}$. Hence $q_{1/2}$ stands for the median operator. We consider ten quantiles $\{\alpha_{i}\}_{i}$ such that $1-\alpha_{i}$ are logarithmically spaced between $10^{-4}$ and $0.4$. Low quantiles are not considered, because, for diffuse observational data, they are contaminated by the noise and the CIB (Ossenkopf-Okada et al., 2016). This logarithmic binning of the high column density values is motivated by the log-normal to power-law behavior of the tail of MCs’ PDFs. See for instance Pouteau et al. (2023) and references therein. Using quantile functions also allows the description to be very robust to outlying pixels, which in practice only drive the last quantile.
@@ 4.2 Two-point based statistics
A very popular way to describe spatial properties of a process is through the PS. It is defined as the Fourier transform of the autocorrelation function, and describes the energy distribution scale by scale in the process studied. In this paper, we consider the isotropic PS, defined as
@@
A power-law behavior $PS[k]\propto k^{-\beta}$ is expected for fields arising in turbulent MHD (Schekochihin, 2022). Thus we adopted a log-log representation to linearize it and for stability purposes (Bruna & Mallat, 2013):
@@
where $x_{\text{apo}}$ is apodized to mitigate non periodic boundary conditions (PBC) as explained in Appendix D. We use six band-pass filters $\{\psi_{i}\}_{1\leq i\leq 6}$ defined as
$\tilde{\psi}_{i}[\@vec{k}]=\mathrm{1}_{\|\@vec{k}\|\in[k_{i},k_{i+1}[},$ (11)
@@ 4.3 Scattering statistics
The wavelet scattering transform (WST) is a set of non-Gaussian descriptors with a hierarchical and multiscale structure (Bruna & Mallat, 2013). It has been shown to be highly effective in describing astrophysical fields (see, for instance, Allys et al., 2019; Saydjari et al., 2021). The usual WST consists in two layers of statistics, the first of which depends on a single scale of length $\simeq 2^{j}$ pixels with orientation $\theta$:
@@
    We note that it computes a $L^{1}$ norm of the filtered field, in contrast with the PS that probes a $L^{2}$ norm.
  • $S_{2}^{Iso1}$ and $S_{2}^{Iso2}$, that measure an isotropic coupling between scales and are obtained by a nonlinear fit of the following model (Allys et al., 2019):
    $\log_{2}\,S_{2}(x)[j_{1},j_{2},\theta_{1},\theta_{2}]=S_{2}^{Iso1}(x)[j_{1},j_{2}]\,+\,\\ S_{2}^{Iso2}(x)[j_{1},j_{2}]\cos\big{[}2(\theta_{2}-\theta_{1})\big{]}.$ (14)
@@
In this paper, we furthermore compress both $S_{2}^{Iso1}$ and $S_{2}^{Iso2}$ coefficients that depend on two scales $(2^{j_{1}},2^{j_{2}})$ by only keeping their $j_{2}-j_{1}$ dependency, that is, a dependency on their ratio, by considering the following average:
@@
In practice, we divided the half-plane $[0,\pi[$ in eight different $\theta$ angles, and we considered four scales999 This choice of $j_{\min}$ is such that, for the finest maps of pixel size $6\arcsec$, the dyadic scale $j_{\min}$ corresponds to $24\arcsec$, which remains above the $18\arcsec$ resolution of the observations. Choosing a larger $j_{\max}$ would allow us to consider larger scales. However, such scales often bring limited additional information due to their large variance, while for computational purposes, a larger $j_{\max}$ leads to a significant loss of the surface used in the image $x$ to estimate the scattering coefficients in the non PBC case. Indeed, to mitigate the latter, convolutions are done with local wavelets and then corrupted edges are cropped with a margin size $\propto 2^{j_{\max}}$. between $j_{\min}=2$ and $j_{\max}=5$. This leads to the three possible values $1\leq\delta\leq 3$ for our reduced $S_{2}$ coefficients. These computations are performed using the pywst [^10] https://github.com/bregaldo/pywst Python package (Regaldo-Saint Blancard et al., 2020).
4.4 Overview
The statistics introduced above are compiled in Table 11. In the following, in addition to the sets of summary statistics defined above, we also consider some aggregated sets of summary statistics that are made from groups of these building blocks. For instance, we refer to $\phi_{\text{Gaussian}}$ statistics to mean the joint set of $\{\phi_{\text{mean}},\phi_{\text{PS}}\}$. These groups are defined in Table 3.
Table 2: Overview of the sets of summary statistics used.
@@ 4.4 Overview
Gaussian description coupling
@@
 [^11] We report their dimension, whether they are sensitive to non-Gaussianity, provide a multiscale description, and probe couplings between scales. Here, “log-Gauss.” reminds us that, even if the statistics at stake are not Gaussian, they remain a very partial diagnostic of non-Gaussianity as these are Gaussian with respect to the logarithm of the field.
Table 3: Aggregated sets of summary statistics used in this work.
@@
final log-Gaussian + RWST 17
5 Toward a low-degeneracy set of statistics
We wanted to construct a set of informative summary statistics for the observational dataset. To do so, we followed a bottom-up approach, starting with the typical low-order statistics and increasingly trying to improve on them by exhibiting and lifting potential degeneracies. The underlying idea is that low-order statistics are good candidates to concentrate most of the informative power into a few coefficients. Such a concentration property is of interest, especially as our low data regime prevents us from learning many features. Another benefit of this approach is to first promote simple statistics, and use more elaborate ones only if they add a significant contribution.
@@ 5.1 Molecular clouds have Gaussian degeneracies
To begin with, we investigate whether Gaussian statistics (i.e., the set made of mean and PS statistics) are degenerate for the observational data. To point out such potential degeneracies, we confront these Gaussian statistics, starting with a low-order non-Gaussian set of statistics, the quantile function QF, using the methodology introduced in Sect. 3.3. As shown in Fig. 6, this first confrontation on observational data already underlines a strong degeneracy level for both Gaussian and QF features, showing that neither statistics is by itself sufficient, according to the compatibility diagnostic we introduced. In particular, the quantile statistics extract a significant amount of information that cannot be efficiently captured by Gaussian statistics.
@@
Figure 6: Confronting Gaussian statistics with QF statistics on observational data, based on the test presented in Fig. 5. Each set of statistics has strong degeneracies lifted by the other set. To investigate the Gaussian confusions, we pick three pairs of $512\times 512$ patches, whose locations on the scatter plot are given by the red stars. These pairs are shown in Fig. 7.
In addition to this dataset-wide diagnostic, we display in Fig. 7 three randomly selected pairs of $512\times 512$ patches degenerate for Gaussian statistics but with increasing QF incompatibility. The locations of these pairs on the scatter plot of Fig. 6 are given there by the red stars. For these three pairs of patches, we show in Fig. 8 the two sets of statistics used in the diagnostic: QF statistics and Gaussian statistics.
@@
Figure 7: Examples of Gaussian confusions. Three pairs of $512\times 512$ patches are chosen, whose locations on the scatter plot of Fig. 6 are given there by the red stars. The column density maps are shown in units of $10^{20}\text{cm}^{-2}$. For each patch, we report: $s$ the subsampling factor from the original 3″/pix map, $d$ and $b$ the approximated distance and Galactic latitude of the cloud. The pixel size (in  mpc) of a patch is thus proportional to $s\times d$. If a pair has patches $(i,j)$ with incompatible pixel sizes, that we define according to the following criterion $\max\{\frac{s_{i}d_{i}}{s_{j}d_{j}},\frac{s_{j}d_{j}}{s_{i}d_{i}}\}\geq 3/2$, we color its labels in red.
@@
Figure 8: Statistics for the examples of Gaussian confusions shown in Fig. 7. In each row, the orange filled line (resp. band) corresponds to the mean (resp. standard deviation) of the statistics computed over the four $256\times 256$ subpatches of the top patch of each pair of Fig. 7, and the corresponding blue lines and areas refer to the bottom patch of the pair. The top row corresponds to the statistics used in the $x$-axis of the scatter plot of Fig. 6, i.e., QF statistics, plotted with ten increasing quantile values, while the bottom row corresponds to the $y$-axis, i.e., Gaussian statistics, plotted starting with mean and followed by the binned PS with six decreasing scales. To better highlight the discrepancies between the two patches of a given pair, we report in second and third rows the offsets of the orange and blue filled lines with respect to their common mean.
Figure 9: Top row: Plots confronting QF statistics with $\{\phi_{\text{mean}},\,\phi_{\text{var}}\}$. Bottom row: Plots confronting QF statistics with $\{\phi_{\text{mean of log}},\,\phi_{\text{var of log}}\}$. These results evidence that taking the logarithm of the map enhances the discriminative power of the mean and variance statistics on both observation and logFBM data (left and middle columns) but not on DTD (right column).
@@
In Fig. 8, the orange filled lines (resp. bands) correspond to the mean (resp. standard deviation) of the statistics computed over the four $256\times 256$ subpatches of the top patch of each pair of Fig. 7, and the corresponding blue lines and areas refer to the bottom patch of the pair. To better highlight the discrepancies between the two patches of a given pair, the offsets of the orange and blue filled lines with respect to their common mean are also shown in the middle rows of Fig. 8.
These results illustrate the ability of QF statistics to distinguish between images with the same Gaussian statistics, which confirms that our compatibility diagnostic works as expected. We emphasize here that the $\phi_{\text{QF}}$ statistics are complementary to the one-point properties probed by Gaussian statistics.
@@
This behavior is not particularly surprising as one-point properties of column density maps of MCs are expected to be at least log-normal, if not with a power-law tail. Hence, estimating such properties directly from the logarithm of the maps, as does $\phi_{\text{QF}}$ but not $\phi_{\text{Gaussian}}$, might enhance the discriminative power. To test this idea, we study in the first column of Fig. 9 the degeneracy level of the set of mean and variance statistics estimated respectively on the raw maps and their logarithm. More precisely, top row plots confront QF statistics with $\{\phi_{\text{mean}},\,\phi_{\text{var}}\}$, while bottom row plots confront QF statistics with $\{\phi_{\text{mean of log}},\,\phi_{\text{var of log}}\}$. We see in this figure that the pairs of observations are clearly better discriminated when using the mean and variance computed on the logarithm of the maps (d) rather than directly (a). This effect is very well reproduced by logFBM data (middle column), but does not hold in general, as shown with the DTD (right column).
@@
Surprisingly, this analysis also shows that, for datasets such as observations or logFBMs, probing the PDF properties through the prism of a mean and a variance is more discriminative than probing its shape. Indeed, in both plots (d) and (e) of Fig. 9, our suitably constructed set $\{\phi_{\text{mean of log}},\,\phi_{\text{var of log}}\}$, of dimension two, performs almost always a better discrimination than the set $\phi_{\text{QF}}$, of dimension ten. This emphasizes the importance of suitably constructed low-dimensional descriptions in such analysis. The effectiveness of such low-dimensional features in the diagnostic we introduced reflects their ability to exploit underlying regularities in the processes at play, enabling data compression with only moderate loss of information. However, when dealing with datasets made of a wide diversity of irregular processes, such as everyday life textures (e.g., DTD), it is hard to identify a compression that does not lead, for some pairs, to poorer performances (c, f).
@@ 5.2 Molecular clouds have log-Gaussian degeneracies
Figure 10: Confrontation of log-Gaussian statistics with RWST statistics on logFBM synthetic data (a), observations (b), simulations (c), and observation-simulation pairs (d). As expected, no log-Gaussian degeneracies are found for the logFBM data (a). However, some are found for the other cases (b,c,d). To investigate the log-Gaussian degeneracies in the observations, six pairs of patches, whose locations on the scatter plot (b) are given by the red stars, are shown in Fig. 19. Only five stars are actually visible, but two of them are overlapped at the horizontal coordinate $x=10$.
@@
The ability of Gaussian one-point statistics to grasp efficiently one-point properties of observations from the logarithm of the column density maps, shown in Fig. 9, suggests to shift toward log-Gaussian statistics, that is, mean and PS estimated on the logarithms of the maps. We thus investigate now whether we may point out some degeneracies of these statistics on observational data. However, if we confront this set with QF statistics, as we did previously for Gaussian statistics, we do not expect to lift significant log-Gaussian degeneracies. Instead, we suggest searching for such degeneracies using a set of higher-order statistics: the RWST. We confront, in Fig. 10, log-Gaussian statistics with the RWST statistics, on three datasets: logFBM in (a), observations in (b) and simulations in (c).
For logFBM data (Fig. 10.a), we see that the RWST diagnostic does not provide any additional information beyond what is already captured by log-Gaussian statistics. Indeed, as expected for such a dataset, log-Gaussian statistics are sufficient (Cover & Thomas, 2006). This result, without being a complete validation, is a clear success of our compatibility diagnostic and supports its relevance. We note that the authors do not include one-point statistics in the set of RWST coefficients, which explains why this descriptor is extremely degenerate for this dataset.
@@
When applying the same analysis on observation data (Fig. 10.b), we see again that the log-Gaussian statistics lift strong RWST degeneracies. However, contrary to logFBM data, we now see that the RWST also lift some degeneracies of the log-Gaussian statistics. To investigate more these degeneracies, we show , in Appendix E, six examples of degenerate pairs in Fig. 19 and the statistics of these patches in Fig. 20. In most of these pairs, the $S_{1}^{Iso}$ coefficients heavily contribute to lift the log-Gaussian degeneracy. In the first and last pairs, $\langle S_{2}^{Iso2}\rangle_{j_{2}-j_{1}}[\delta]$ and respectively $\langle S_{2}^{Iso1}\rangle_{j_{2}-j_{1}}[\delta]$ are also very discriminative.
@@
Finally, simulations yield yet another result (Fig. 10. c). Although this case leads to the same qualitative conclusions as observations, namely underlining the insufficiency of log-Gaussian statistics, it differs from it quantitatively. This shows that the distribution of simulations in the space of summary statistics is different from that of observations. In particular, the scatter plots (b) and (c) of Fig. 10 show that the structures of these distributions, that is, the dependency between the statistics for each of them, are not the same. This supports the caveat of simulation-based inference according to which observations and simulations of such ISM processes have different statistical structures, and motivates the observation-based approach of this paper.
To complement the previous results, we show the scatter plot obtained by directly comparing patches from observations and simulations (Fig. 10.d). This plot reveals even greater degeneracies than those observed previously, further highlighting the difference between observations and simulations. We note that this also leaves open the question of whether the set of log-Gaussian + RWST statistics also has degeneracies of its own to make such a comparison. This could be further investigated by confronting it with other sets of statistics using the diagnostic introduced in this paper.